solve by using elimination and substitution method:
3x+4y=6 and 2x+3y=5

Answer by checkley77(12844)

3x+4y=6 multiply by 2
2x+3y=5 multiply by -3 & add.
6x+8x=12
-6x-9x=-15
--------------
-x=-3
y=3 ans.
3x+4*3=6
3x+12=6
3x=6-12
3x=-6
x=-6/3
x=-2 ans.
------------------------------------------------
3x=-4y+6
x=-4y/3+6/3
x=-4y/3+2
2(-4y/3+2)+3y=5
-8y/3+4+3y=5
-8y/3+3y=5-4
(-8y+3y*3)/3=1
-8y+9y=3*1
y=3 ans.
3x=-4*3+6
3x=-12+6
3x=6
x=6/3
x=2 ans.